### v2

# On synchronized coded systems

Abstract.
We introduce a class of codes with overlapping code words, that we call SPO-codes. The SPO-codes are related to the Markov codes that were introduced in: G. Keller, J. Combinatorial Theory 56, (1991), pp. 75–83. The process of generating a coded system from a code extends to SPO-codes. We describe a family of intrinsically ergodic synchronized SPO-coded systems that is closed under topological conjugacy. We construct synchronized subshifts with salient structural features by means of SPO-codes. We construct SPO-coded systems that are not semisynchronized.
footnotetext: MSC2020: Primary 37B10, Secondary 68Q45.
1. Introduction
Let $\Sigma$ be a finite alphabet, and let $S_{\Sigma}$ denote the left shift on $\Sigma^{\mathbb{Z}}$. Given an $S_{\Sigma}$-invariant subset $X$ of $\Sigma^{\mathbb{Z}}$ the restriction $S_{\Sigma}$ of $S_{\Sigma}$ to $X$ is called a subshift. For basic notions of the theory of subshifts we refer to [Ki] or [LM]. A word is said to be admissible for a subshift $X\subset\Sigma^{\mathbb{Z}}$ if it appears on a point of $X$. We denote the set of admissible words of the subshift $X$ by $\mathcal{L}(X)$.
Coded systems were introduced in [BH]. Given the finite alphabet $\Sigma$, a set of words in the symbols of $\Sigma$ is called a code. The set of points in $\Sigma^{\mathbb{Z}}$ that carry a bi-infinite concatenation of words in $\mathcal{C}$ is called the concatenation set of the code. A subshift that is the closure of the concatenation set of a code is called a coded system.
Before indicating the content of the paper we introduce terminology and notation. Given a subshift $X\subset\Sigma^{\mathbb{Z}}$, and $i^{-},i^{+}\in\mathbb{Z},i^{-}\leq i^{+}$ we use the notation
$x_{[i^{-},i^{+}]}=(x_{i})_{i^{-}\leq i\leq i^{+}},\quad\quad\quad\quad(x\in X),$
and the notation
$X_{[i^{-},i^{+}]}=\{x_{[i^{-},i^{+}]}:x\in X\}.$
We use similar notations in the case that indices range in left infinite or right infinite intervals. We identify the blocks in $X_{[1,l)},l\in\mathbb{N},$ with the words that they carry. We use the same notation for a finite or a left- or right-infinite block and for the finite or a left- or right-infinite word that the block carries. Concatenations of left-infinite, finite and right-infinite blocks and words we denote by juxtaposition.
Given a subshift we use notations like $x^{\langle-\rangle}$ $(x^{\langle+\rangle})$ for the points in $X_{(-\infty,i]}(X_{[i,\infty)}),$ $i\in\mathbb{Z}$. The right context of the finite block resp. finite word $a$ is denoted by
$\Gamma^{\langle+\rangle}_{\infty}(a)=\{x^{\langle+\rangle}\in X_{(j,\infty)}:ax^{\langle+\rangle}\in X_{[i,\infty)}\},\qquad(a\in x_{[i,j]},i,j\in\mathbb{Z},i\leq j),$
$\Gamma^{\langle+\rangle}(a)=\{b\in\mathcal{L}(X):ab\in\mathcal{L}(X)\},\qquad\qquad\qquad(a\in\mathcal{L}(X)).$
The notation $\Gamma^{\langle-\rangle}$ has the symmetric meaning. We also set
$\omega^{\langle+\rangle}(a)=\bigcap_{x^{\langle-\rangle}\in\Gamma_{\infty}^{\langle-\rangle}(a)}\{x^{\langle+\rangle}\in\Gamma_{\infty}^{\langle+\rangle}$ $(a):x^{\langle-\rangle}ax^{\langle+\rangle}\in X\},$
$(a\in x_{[i,j]},\ i,j\in\mathbb{Z},i\leq j).$
The notation $\omega^{\langle-\rangle}$ has the symmetric meaning.
Given a subshift $X\subset\Sigma^{\mathbb{Z}}$ a point $x^{\langle-\rangle}\in X_{(-\infty,i]}(x^{\langle+\rangle}\in X_{(-\infty,i]})$$i\in\mathbb{Z},$ is said to be transitive if for all $j<i$ ($j>i$) all words in $\mathcal{L}(X)$ appear in $x^{\langle-\rangle}_{(-\infty,j]}$($x^{\langle+\rangle}_{[j,\infty)}$).
Synchronized subshifts were introduced in [BH]. A word $c\in\mathcal{L}(X)$ is synchronizing if one has for $u^{\langle-\rangle}\in\Gamma^{\langle-\rangle}(c)$ and $u^{\langle+\rangle}\in\Gamma^{\langle+\rangle}(c)$ that $u^{\langle-\rangle}cu^{\langle+\rangle}\in X$. A topologically transitive subshift is said to be synchronized if it has a synchronizing word. The set of synchronizing words of a synchronized subshift $X$ we denote by $\mathcal{L}_{synchro}(X)$. A word in $\mathcal{L}(X)$ that has a subword in $\mathcal{L}_{synchro}(X)$ is in $\mathcal{L}_{synchro}(X)$.
Semisynchronizing subshifts were introduced in [Kr1]. A word $a$ is said to be semisynchronizing if $\omega^{-}(a)$ contains a transitive point. A subshift is called semisynchronizing if it has a semisynchronizing word.
Hyposynchronized subshifts were introduced in [Kr2]. A word $a\in\mathcal{L}(X)$ is hyposynchronizing iif $\omega^{+}(a)$ contains a transitive point, and if there exists a point
$x^{\langle-\rangle}\in\Gamma^{\langle-\rangle}(a)$
such that
$\omega^{\langle+\rangle}(x^{\langle-\rangle}_{(-k,0]}a)=\omega^{\langle+\rangle}(a),\qquad(k\in\mathbb{N}).$
A subshift is hyposynchronized if it has a hyposynchronizing word. The inverse of a hyposynchronized subshift is semisynchronized.
For a subshift $X\subset\Sigma^{\mathbb{Z}}$ we denote by $\mathcal{L}_{Markov}(X)$ the set of words $a\in\mathcal{L}(X)$ such that
$\{\Gamma^{+}(ba):b\in\Gamma^{-}(a)\}=\infty.$
The subshift that has $\mathcal{L}_{Markov}(X)$ as its set of admissibble words was introduced in [T] and named the Markov boundary of $X$. We denote the Markov boundary of $X$ by $\partial_{Markov}(X)$.
The content of this paper is as follows. Section 2 contains preliminaries. In particular it contains a lemma on synchronizing words and a lemma on semisynchronizing words.
[…]
In Section 4 we show that to a synchronized subshift $X\subset\Sigma^{\mathbb{Z}}$ there is associated a PSO-code $\mathcal{C}(X)$. This SPO-code $\mathcal{C}(X)$ is canonical in the sense that its concatenation set is invariantly attached to $\mathcal{C}(X)$. As an application of the condition of Section 3 we give a sufficient condition on a synchronizing subshift to be intrinsically ergodic. In the situation that we consider the set of synchronizing subshifts that satisfy this condition is closed under topological conjugacy.
In Section 5 we construct by means of SPO-codes synchronized shifts with salient structural properties. We construct SPO-codes $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ consists of a finite number of orbits, and SPO-codes $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ consists of a countable infinity of orbits.
In Section 6 we describe coded systems that are non-synchronized in the sense that they are neither semisynchronized nor hyposynchronized. These coded systems and their variations can be considered as structurally distant from the Dyck shifts.
2. Preliminaries
We introduce more notation and terminology. Consider a finite alphabet $\Sigma$. For $\sigma\in\Sigma$ we denote the word $a$ of length $n\in\mathbb{N}$ that is given by $a_{i}=\sigma,1\leq i\leq n,$ by $\sigma^{n}$. The length of a finite word $a$ we denote by $\ell(a)$. Beyond subshifts we consider more generally restriction $S_{X}$ of the shift $S_{\Sigma}$ to shift invariant Borel subsets of $\Sigma^{\mathbb{Z}}$. We call such restrictions Borel shifts. In the case that a Borel shift possesses invariant probability measures its topological entropy $h(X)$ is defined as the supremum of the measure theoretic entropies of its invariant probability measures.
We recall that, given finite alphabets $\Sigma$ and $\bar{\Sigma}$, subshifts $X\subset\Sigma^{\mathbb{Z}},\bar{X}\subset\bar{\Sigma}^{\mathbb{Z}}$, and a shift commuting continuous map
$\varphi:X\to\bar{X},$
one has some $L\in\mathbb{N}$ such that [L.L][-L.L] is the coding window of a block map
$\Phi:\Sigma^{[-L.L]}\to\bar{\Sigma}$
that implements $\varphi$, in the sense that
$\varphi((x_{i})_{i\in\mathbb{Z}})=\Phi(x_{[i-L.L+i]})_{i\in\mathbb{Z}},\quad\qquad\qquad((x_{i})_{i\in\mathbb{Z}}\in X).$
Given a coding window [L,L][-L,L] and a bock map
$\Phi:\Sigma^{[-L.L]}\to\bar{\Sigma}$
we use the notation
$\Phi(x^{\langle-\rangle})=(x^{\langle-\rangle}_{[i-L,i+L]})_{-\infty<i<j-L},\qquad(x^{\langle-\rangle}\in X_{(-\infty,j},j\in\mathbb{Z}).$
That the family of synchronized shifts is closed under topological conjugacy follows from the following lemma that we include for completeness.
Lemma 1.
Let $X\subset\Sigma^{\mathbb{Z}},\bar{X}\subset\bar{\Sigma}^{\mathbb{Z}}$ be topologically conjugate subshifts and let
$\varphi:X\to\bar{X}$
be a topological conjugacy . Let [L,L][-L,L] be a coding window for $\varphi$ and for $\varphi^{-}$. Let $x\in X$ such that
(1) $x_{[-L,L]}\in\mathcal{L}_{synchro},$
and set $\bar{x}=\varphi(x)$. Then
(2) $\bar{x}_{[-2L,2L]}\in\mathcal{L}_{synchro}(\bar{X}).$
Proof.   Let $\varphi$ be implemented by a block map
$\Phi:\Sigma^{[-L.L]}\to\bar{\Sigma}$
and let $\varphi^{-}$ be implemented by a block map
$\bar{\Phi}:\bar{\Sigma}^{[-L.L]}\to\Sigma.$
Let
$\bar{y}^{\langle-\rangle}\in\Gamma^{\langle-\rangle}(\bar{x}_{[-2L,2L]}),$
and let $y^{\langle-\rangle}\in X_{(-\infty,-L)}$ be given by
$y^{\langle-\rangle}=\bar{\Phi}(\bar{y}^{\langle-\rangle},\bar{x}_{[-2L,0)}).$
One has that
$\bar{\Phi}(\bar{y}^{\langle-\rangle},\bar{x}_{[-2L,2L)})=(y^{\langle-\rangle},x_{[-L,L)}).$
Therefore
$y^{\langle-\rangle}\in\Gamma^{\langle-\rangle}(x_{[-L,L)}).$
Symmetrically, let
$\bar{y}^{\langle+\rangle}\in\Gamma^{\langle+\rangle}(\bar{x}_{[-2L,2L]}),$
and set
$y^{\langle+\rangle}=\bar{\Phi}(\bar{y}^{\langle+\rangle},\bar{x}_{,2L0)}),$
and have that
$y^{\langle+\rangle}\in\Gamma^{\langle+\rangle}(x_{[-L,L)}).$
It follows by (1) that
$(y^{\langle-\rangle},x_{[-L,L)},y^{\langle+\rangle})\in X.$
One checks that
$\varphi(y^{\langle-\rangle},x_{[-L,L)},y^{\langle+\rangle}))=(\bar{y}^{\langle-\rangle},x_{[-2L,2L)},\bar{y}^{\langle+\rangle}\in\bar{X}.$
This proves (2). $\Box$
For semisynchronizing words there is the following lemma.
Lemma 2.
Let $X\subset\Sigma^{\mathbb{Z}}$ be a semisynchronizing subshift and let $a\in\mathcal{L}(X)$ be a semisynchronizing word of $X$. Let
(3) $c\in\Gamma^{\langle-\rangle}(a).$
Then the word $ac$ is semisynchronizing.
Proof.   Let
(4) $x^{\langle-\rangle}\in\omega^{\langle-\rangle}(a)$
be transitive. One shows that
(5) $x^{\langle-\rangle}\in\omega^{\langle-\rangle}(ac).$
For this one has to show for
(6) $y^{\langle+\rangle}\in\Gamma_{\infty}^{\langle+\rangle}(ac),$
that
(7) $(x^{\langle-\rangle},ac,y^{\langle+\rangle})\in X.$
One has from (3) and (6) that
$cy^{\langle+\rangle}\in\Gamma_{\infty}^{\langle+\rangle}(a).$
From this one sees that (7) follows from (4). $\Box$
3. Markov codes and SPO-codes
3.1 $\mathbb{Markov\ codes.}$ Let there be given a finite alphabet $\Sigma$. A Markov code 11 We consider codes that are Markov codes in the sense of Keller [Ke] is given by a code $\mathcal{D}\subset\Sigma^{\mathbb{Z}}$ as state space and an irreducible 0 - 1 transition2 2For the countable state Bernoulli case see [KSW]. matrix $T_{\mathcal{D}}(d,d^{\prime}),d,d^{\prime}\in\mathcal{D}$. The concatenation set $C(\mathcal{D})$ of a Markov code $\mathcal{D}$ is the set of points $x\in\Sigma^{\mathbb{Z}}$ such that there exists a sequence of indices $i_{k},k\in\mathbb{Z},$
$i_{k-1}<i_{k},\qquad(k\in\mathbb{Z},)$
such that
(8) $x_{(i_{k-1},i_{k}]}\in\mathcal{D},$
$T_{\mathcal{D}}(x_{(i_{k-1},i_{k}]},x_{(i_{k},i_{k+1}]})=1,\qquad\qquad(k\in\mathbb{Z}).$
[…]
$i_{0}\leq 0<i_{1},$
[…]
3.2. $\mathbb{SPO-codes.}$ For a finite alphabet $\Sigma$ let there be given a bifix code $\mathcal{F}\subset\Sigma^{+}$. Denote by $\mathcal{C}_{\mathcal{F}}$ the set of words $c\in\Sigma^{+}$ that have a proper prefix $f^{-}(c)\in\mathcal{F}$ and also a proper suffix $f^{+}(c)\in\mathcal{F}$. For a word $c\in\mathcal{F}$ we denote by $\mathring{c}$ the word that is obtained by removing from $c$ the suffix $f^{+}(c)$.
For $a,b\in\mathcal{C}_{\mathcal{F}}$ we define the concatenation with overlap $a\circledast b$ by setting
[…]
The word $a\circledast b$ is in $\mathcal{C}_{\mathcal{F}}$:
$f^{-}(a\circledast b)=f^{-}(a),\quad f^{+}(a\circledast b)=f^{+}(b).$
For $Q>2$ and for
$c_{q}\in\mathcal{C}_{\mathcal{F}},\quad\quad 1\leq q\leq Q,$
we set
$\prod^{\circledast}_{1\leq q\leq Q}c_{q}=(\prod^{\circledast}_{1\leq q<Q}c_{q})\circledast c_{Q}.$
We define an SPO-code as a code $\mathcal{C}$ that is contained in $\mathcal{C}_{\mathcal{F}}$ for some bifix code $\mathcal{F}$. The concatenation set $C(\mathcal{C})$ of an SPO-code $\mathcal{C}$ is the set of $x\in\Sigma^{\mathbb{Z}}$ such that there exists a sequence of indices $(j_{j},i_{k})\in\mathbb{Z}^{2}$ such that
(9) $i_{k-1}<j_{k}\leq i_{k},\qquad(k\in\mathbb{Z}),$
and such that
(10) $x_{(i_{k-1}.i_{k}]}\in\mathcal{C},\ \ x_{(j_{k}.i_{k}]}=f^{+}(x_{(i_{k-1}.i_{k}]})=f^{-}(x_{(i_{k},i_{k+1}]}),\qquad(k\in\mathbb{Z}).$
An SPO-code $\mathcal{C}$ is said to be unambiguous if for all points $x$ in its concatenation set the sequence. of indices $(j_{j},i_{k})\in\mathbb{Z}^{2},$ that is assumed normalized,
$i_{0}\leq 0<i_{1},$
such that (8) and (9) hold, is unique. The closure of $C(\mathcal{C})$, to be denoted by $X(\mathcal{C})$, is the subshift that is generated by $\mathcal{C}$.
Given a bifix code $\mathcal{F}$ we set
$\mathcal{C}_{\mathcal{F}}^{\bullet}=\{c\in\mathcal{C}_{\mathcal{F}}^{\circ}:\ell(c)\leq\ell(f^{-}(c))+\ell(f^{+}(c))\}.$
Proposition 3.
Given a bifix code $\mathcal{F}$, let $\mathcal{C}\subset\mathcal{C}_{\mathcal{F}}^{\circ}$ be an SPO-code with irreducible transition matrix, such that
$\mathcal{C}\cap\mathcal{C}_{\mathcal{F}}^{\bullet}\neq\emptyset,$
and
(11) $C(\mathcal{C}))>h(X(\mathcal{C})\setminus C(\mathcal{C})),$
and such that $X(\mathcal{C})$ has a measure of maximal entropy of full support. Then $X(\mathcal{C})$ is intrinsically ergodic.
Proof.   Given a point $x\in C(\mathcal{C})$ denote by
$(j_{k}(x),i_{k}(x))_{k\in\mathbb{Z}}$
the unique sequence
$(j_{k}(x),i_{k}(x))_{k\in\mathbb{Z}}\in\mathbb{Z}^{2},$
such that
$i_{k-1}<j_{k}\leq i_{k},\qquad(k\in\mathbb{Z}),$
with the normalization
$i_{0}\leq 0<i_{1},$
and such that
$x_{(i_{k-1},i_{k}]}\in\mathcal{C},\quad x_{({j_{k}},i_{k}]}=f^{+}(x_{(i_{k-1},i_{k}]}),\qquad(k\in\mathbb{Z}).$
We set
$\mathcal{K}(x)=\{k\in\mathbb{Z}:x_{(i_{k-1},i_{k}]}\in\mathcal{C}_{\mathcal{F}}^{\bullet}\},\qquad\qquad(x\in C(\mathcal{C})).$
From the unambiguous SPO-code $\mathcal{C}$ with irreducible transition matrix $T_{C}$ we obtain an unambiguous SPO-code $\widehat{\mathcal{C}}$ with irreducible transition matrix $T_{\widehat{C}}$ by setting
(12) $\widehat{\mathcal{C}}=\{(\prod^{\circledast}_{1\leq q<Q}c_{q})\circledast c:c\in\mathcal{C}_{\mathcal{F}}^{\bullet},c_{q}\in\mathcal{C}\setminus\mathcal{C}_{\mathcal{F}}^{\bullet}\,1\leq q<Q,Q\in\mathbb{N}\}.$
[…]
$(\widehat{j}_{k}(x),\widehat{i}_{k}(x))_{k\in\mathbb{Z}}$
the unique sequence
$(\widehat{j}_{k}(x),\widehat{i}_{k}(x))_{k\in\mathbb{Z}}\in\mathbb{Z}^{2},$
such that
$\widehat{i}_{k-1}<\widehat{j}_{k}\leq\widehat{j}_{k},\qquad(k\in\mathbb{Z}),$
that is assumed normalized,
$\widehat{i}_{0}\leq 0<\widehat{i}_{1},$
and such that
$x_{(\widehat{i}_{k-1},\widehat{i}_{k}]}\in\widehat{\mathcal{C}},\quad x_{({j\widehat{j}_{k}},\widehat{i}_{k}]}=f^{+}(x_{(\widehat{i}_{k-1},\widehat{i}_{k}]}),\qquad(k\in\mathbb{Z}).$
We set
$\mathcal{D}_{\mathcal{C}}=\{\mathring{c}:c\in\widehat{\mathcal{C}}\}$.
By construction (see (12)) the words $d\in\mathcal{D}_{\mathcal{C}}$ have a prefix $f^{-}(d)\in\mathcal{F}$. We have obtained an unambiguous Markov code with state space $\mathcal{D}_{\mathcal{C}}$ and transition matrix $T_{\mathcal{D}_{\mathcal{C}}}$ that is given by
[…]
The concatenation set of the Markov code $\mathcal{D}_{\mathcal{C}}$ coincides with the concatenation set of $\widehat{\mathcal{C}}$. From the Markov code $\mathcal{D}_{\mathcal{C}}$ we obtain a topological Markov shift $Y(\mathcal{C})$ with state space $\mathcal{S}$ given by
$\mathcal{S}=\{(d,l):d\in\mathcal{D}_{\mathcal{C}},1\leq l\leq\ell(d)\},$
and transition matrix $T_{\mathcal{S}}$ given by
$T_{\mathcal{S}}((d,l),(.d,l+1))=1, \quad d\in\mathcal{D}_{\mathcal{C}},1\leq l<\ell(d),$
and
$T_{\mathcal{S}}((d,\ell(d)),(d^{\prime},1)=\begin{cases}1,\quad\text{if $T_{\mathcal{D}_{\mathcal{C}}}(d,d^{\prime})=1$},\\ 0,\quad\text{if $T_{\mathcal{D}_{\mathcal{C}}}(d,d^{\prime})=0$}.\end{cases}$
As a dynamical system $C(\mathcal{D}_{\mathcal{C}})$ is Borel conjugate to $Y(\mathcal{C})$. A Borel conjugacy
$\varphi:C(\mathcal{D}_{\mathcal{C}})\to Y(\mathcal{C})$
maps a point $x\in C(\mathcal{D}_{\mathcal{C}})$ to the point $(s_{i})_{i\in\mathbb{Z}}$ that is determined by
$s_{0}(x)=(x_{(\widehat{i}_{0},\widehat{i}_{1}]},-\widehat{i}_{0}).$
One has that
$C(\widehat{\mathcal{C}})\subset C(\mathcal{C}).$
[…]
In imposing the hypothesis (11) we have followed the approach of [KSW, Theorem 3].
4. Synchronization
4.1. $\mathbb{The\ SP0-code\ of\ a\ synchronized\ subshift}.$ Given a synchronized subshift $X\subset\Sigma^{\mathbb{Z}}$ we set
$\mathcal{J}_{i}(x)=$ $\{j\in(-\infty,i]:x_{[j,i]}\in\mathcal{L}_{synchro}(X)\},\qquad(x\in X,i\in\mathbb{Z}),$
$B^{\circ}(X)=\bigcap_{i\in\mathbb{Z}}\{x\in X:\mathcal{J}_{i}(x)\neq\emptyset\},$
$J_{i}(x)=$ $\max\thinspace(\mathcal{J}_{i}(x)),\qquad\qquad\ \ \ \qquad(x\in X,i\in\mathbb{Z},x\in B^{\circ}(X)).$
Lemma 4.
For a synchronized subshift $X\subset\Sigma^{\mathbb{Z}}$ one has that
$J_{i+1}(x)\geq J_{i}(x),\qquad\qquad(x\in X,i\in\mathbb{Z},x\in B(X)).$
Proof.   The lemma follows from
$x_{[J_{i},i+1]}\in\mathcal{L}_{synchro}(X),\qquad\qquad(x\in X,i\in\mathbb{Z}).\qed$
$\Box$ We denote
$B(X)=\{x\in B^{\circ}(X):\lim_{i\to\infty}J_{i}(x)=\infty\}.$
One notes that the set $B(X)$ is Borel.
Lemma 5.
For an synchronized subshift $X$ that has an measure of full support $B(X)$ is not empty.
Proof.   A measure of full support of the synchronized subshift $X$ assigns positive measure to $B(X)$. $\Box$ The next lemma shows that the set $B(X)$ is invariantly attached to $X$.
Lemma 6.
Let $X\subset\Sigma^{\mathbb{Z}},\bar{X}\subset\bar{\Sigma}^{\mathbb{Z}}$ be topologically conjugate subshifts and let
$\varphi:\bar{X}\to X$
be a topological conjugacy . Let [L,L][-L,L] be a coding window for $\varphi$ and for $\varphi^{-}$.Let $x\in X$ such that
$\lim_{i\to\infty}J_{i}(x)=\infty,$
and set $\bar{x}=\varphi^{-1}$. Then
$\lim_{i\to\infty}J_{i}(\bar{x})=\infty.$
Proof.   Let $\bar{k}\in\mathbb{Z}$, and let $i\in\mathbb{Z}$ such that
$J_{i}(x)>\bar{k}+L.$
then
$x_{(J_{i}(x),i]}\in\mathcal{L}_{synchro}(X).$
By Lemma 1 then
$\bar{x}_{(J_{i}(\bar{x})-L,i+L]}\in\mathcal{L}_{synchro}(\bar{X}).$
This means that
$J_{i+L}(\bar{x})\geq\bar{k}.$
$\Box$
For $x\in B(X)$ we denote by $\mathcal{I}(x)$ the set of $i\in\mathbb{Z}$ such that
$J_{i}(x)>J_{i-1}(x).$
and we enumerate $\mathcal{I}(x)$,
(13) $\mathcal{I}(x)=\{I_{i}(x):i\in\mathbb{Z}\},$
with the normalization
(14) $I_{0}(x)\leq 0<I_{1}(x),\qquad(x\in B(X)).$
We associate to the synchronizing subshift $X\subset\Sigma^{\mathbb{Z}}$ an SPO-code $\mathcal{C}(X)$. The prefixes (suffixes) of the code $\mathcal{C}(X)$ arise as the words in
$\{x_{(J_{i}(x),i]}:i\in\mathcal{I}(x),x\in B(X)\},$
and the words in $\mathcal{C}(X)$ arise as the words in
$\{x_{(I_{i}(x),I_{i+1}(x)]}:i\in\mathbb{Z},x\in B(X)\}.$
The set $\mathcal{F}_{\mathcal{C}(X)}$ of prefixes (suffixes) of the code $\mathcal{C}(x)$ contains precisely the synchronizing words of $X$ that have no proper subword that is synchronizing. The words in the code $\mathcal{C}(X)$ are precisely the synchronizing words of $X$ that have a proper prefix and a proper suffix in $\mathcal{F}_{\mathcal{C}(X)}$ and that have no other subword that is synchronizing.
The set $B(X)$ is equal to $C(\mathcal{C}(X))$. As the indices in (4.4), (4.5) are uniquely determined by the point $x$, the SPO-code $\mathcal{C}(X)$ is unambiguous.
We formulate a Condition (H) on a synchronizing subshift $X$:
(H) $\sup_{c\in\mathcal{C}(X)}\{\ell(c)-\ell(f^{-}(c))-\ell(f^{+}(c))\}=\infty.$
Lemma 7.
The condition H is an invariant of topological conjugacy of synchronizing shifts.
Proof.   Apply Lemma 1. $\Box$
Theorem 8.
Let $X$ be a topologically transitive synchronized subshift that has a measure of maximal entropy of full support, that satisfies Condition H, and is such that
$h(B(X))>h(X\setminus B(X)).$
Then $X$ is intrinsically ergodic.
Proof.   Apply Proposition 3. $\Box$
5. A class of synchronized shifts
We give an example such that $\partial_{Markov}(X(\mathcal{C}))$ contains finitely many orbits and an example of an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ contains a countable infinity of orbits. The SPO-coded shifts of the examples satisfy Condition H, and are therefore intrinsically ergodic by Theorem 8.
For the construction of the SPO-codes we use the alphabet $\{\delta,\gamma\}$ and a finite alphabet $\Sigma$ that is disjoint from $\{\delta,\gamma\}$. The words $\gamma\delta^{n}\gamma,n\in\mathbb{N},$ will serve as the distinguished suffixes and prefixes of the SPO-codes.
5.1. $\mathbb{Example\ 1.}$ We construct an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ is a finite set of orbits. For the construction we choose a finite set $P$ of periodic points of $(\Sigma^{\mathbb{Z}},S_{\Sigma^{\mathbb{Z}}})$.
Set
$R=\prod_{p\in P}\text{period of}(p).$
The code is given by
$\mathcal{C}=$ $\{\gamma\delta\gamma p_{[0,2qR)}\gamma\delta\gamma:p\in P,q>1\}\cup$
$\{\gamma\delta^{n}\gamma p_{[0,2qR)}\gamma\delta^{n+1}\gamma:p\in P,n\in\mathbb{N},q\geq n\}\cup$
$\{\gamma\delta^{n}\gamma p_{[0,2qR)}\delta^{n-1}:p\in P,n>1,q\geq n\}.$
One has that
$\Gamma(\delta^{n+1}\gamma p_{[0,R)})\begin{cases}\ni p_{[0,nR)}\gamma,\\ \not\ni p_{[0,(n+1)R)}\gamma,\end{cases}\qquad\qquad(p\in P,n\in\mathbb{N}).$
It follows that $\partial_{Markov}(X(\mathcal{C}))$ contains the points in $P$, the fixed point $(x_{i})_{i\in\mathbb{Z}}$ that is given by $x_{i}=\delta,i\in\mathbb{Z},$ and the orbits that contain the points $x^{p,-},x^{(p,+)},p\in P,$ that are given by,
$x^{p,-}_{(-\infty,0)}=p_{(-\infty,0)},\qquad\qquad x^{p,+}_{[0,\infty)}=p_{[0,\infty)},$
$\qquad x^{p,-}_{i}=\begin{cases}\gamma,\text{if $i=1,$}\\ \delta,\text{if $i>1,$}\end{cases}\quad\qquad\qquad x^{p,+}_{i}=\begin{cases}\gamma,\text{if $i=1,$}\\ \delta,\text{if $i>1,$}\end{cases}$
and no other points.
5.2. $\mathbb{Example\ 2.}$ We construct an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ is countable infinity of orbits. We denote by $p^{(k)}$ the point of period $k$ in $(\{0,1\}^{\mathbb{Z}},S_{\{0,1\}^{\mathbb{Z}}})$ that is given by
$p^{(k)}_{i}=\begin{cases}0,\text{if $0\leq i<k-1$,}\\ 1,\text{if $i=k-1$.}\end{cases}$
Set
$R_{k}=1+\tfrac{1}{2}{k(k-1)},\qquad(k\in\mathbb{N}).$
The code is given by
$\mathcal{C}=$ $\{\gamma\delta^{k}\gamma p^{(1)}_{[0,2k)}\gamma\delta^{(R_{k})}\gamma:k\in\mathbb{N}\}\cup$
$\{\gamma\delta^{k}\gamma p^{(m)}_{[0,2mk)}\gamma\delta^{k+m-1}\gamma:1<m\leq k,k\in\mathbb{N}\}.$
Consider the words
$b^{-}(q,k,m)=\delta^{q}\gamma p^{(k)}_{[0,2mk)},\qquad(q,k,m\in\mathbb{N}),$
that are in $\mathcal{L}(\mathcal{C}(X))$. Set
$K^{-}_{\circ}=\begin{cases}k-q,\text{if $q<k,$}\\ 0,\text{if $q=k$.}\end{cases}$
One has that
$\Gamma^{+}(\gamma\delta^{K}b^{-}(m,k,q))\begin{cases}\ni p^{(m)}_{[0,2mk)}\gamma,\\ \not\ni p_{[0,2m(K+1)}\gamma,\end{cases}\qquad\qquad(K\geq K_{\circ}).$
This implies that $b^{-}(q,k,m)\in\mathcal{L}_{Markov}(\mathcal{C}(X))$.
Consider the words
$b^{+}(q,k,m)=\delta^{q}\gamma p^{(k)}_{[0,2km)},\qquad(q,k,m\in\mathbb{N}),$
that are in $\mathcal{L}(\mathcal{C}(X))$. Set
$M^{-}_{\circ}=\begin{cases}q-m,\text{if $q>m,$}\\ 0,\text{if $q\leq m$.}\end{cases}$
One has that
$\Gamma^{+}(\gamma p^{(k)}_{[0,2kM)}b^{+}(q,k,m))\begin{cases}\ni\delta^{R_{\max\{q,m\}+M-M_{\circ}}+\max\{q,m\}+M-M_{\circ}-1-q}\gamma,\\ \not\ni\delta^{R_{\max\{q,m\}+M-M_{\circ}}+\max\{q,m\}+M-M_{\circ}-q},\end{cases}(M\geq M_{\circ}).$
This implies that $b^{+}(q,k,m)\in\mathcal{L}_{Markov}(\mathcal{C}(X))$.
It follows that $\partial_{Markov}(X(\mathcal{C}))$ contains the points $p^{(k)},k\in\mathbb{N}$, the the fixed point $(x_{i})_{i\in\mathbb{Z}}$ that is given by $x_{i}=\delta,i\in\mathbb{Z},$ the fixed point $(y_{i})_{i\in\mathbb{Z}}$ that is given by $y_{i}=0,i\in\mathbb{Z},$ and the orbits that contain the points $x^{k,-},x^{(k,+)},k\in\mathbb{N},$ that are given by,
$x^{(k,-)}_{(-\infty,0)}=p^{(k)}_{(-\infty,0)},\qquad\qquad x^{p,+}_{[0,\infty)}=p^{(k)}_{[0,\infty)},$
$\qquad x^{p,-}_{i}=\begin{cases}\gamma,\text{if $i=1,$}\\ \delta,\text{if $i>1,$}\end{cases}\quad\qquad\qquad x^{p,+}_{i}=\begin{cases}\gamma,\text{if $i=1,$}\\ \delta,\text{if $i>1,$}\end{cases}$
and no other points.
6. A class of coded systems
For the construction of examples of subshifts that are not semisynchronizing we use the alphabet
$\Sigma=\{-1,0,1\}.$
For $k\in\mathbb{Z}_{+},\alpha\in\{-1,1\}$, we denote by $g^{\langle-,k,\alpha\rangle}$ the word of length $k+1$ that is given by
$g^{\langle-,k,\alpha\rangle}_{i}=\begin{cases}0,\ \text{if $1\leq i\leq k,$}\\ \alpha,\ \text{if $i=k+1,$}\end{cases}$
and we denote by $g^{\langle+,k,\alpha\rangle}$ the word of length $k+1$ that is given by
$g^{\langle-,k,\alpha\rangle}_{i}=\begin{cases}\alpha,\ \text{if $i=1,$}\\ 0,\ \text{if $1<i\leq k+1.$}\end{cases}$
For $k\in\mathbb{N},\alpha\in\{-1,1\}$ we denote by $c^{\langle k,\alpha\rangle}$, the word of length $k+2$ that is given by
$c^{\langle k,\alpha\rangle}_{i}=\begin{cases}\alpha,\ \text{if $i=1.$}\\ 0,\ \text{if $1<i\leq k+1.$}\\ \alpha\ \text{if $i=k+2.$}\end{cases}$
Note that
$\alpha$ $=g^{\langle-,0,\alpha\rangle}=g^{\langle+,0,\alpha\rangle},\qquad\quad(\alpha\in\{-1,1\}).$
$c^{\langle k,\alpha\rangle}$ $=g^{\langle+,k^{+},\alpha\rangle}g^{\langle-,k^{-},\alpha\rangle},\qquad\ (\alpha\in\{-1,1\},k^{+},k^{-}\in\mathbb{Z}_{+},k^{+}+k^{-}>0).$
We denote the coded system that is generated by the code
$\mathcal{C}=\{c^{\langle k,\alpha\rangle}:k\in\mathbb{N},\alpha\in\{-1,1\}\}$
by $X_{\mathcal{C}}$. We set
$\mathcal{D}^{\langle 1\rangle}=\{\alpha 0\alpha:\alpha\in\{-1,1\}\},$
$\mathcal{D}^{\langle m+1\rangle}=\{\alpha(\bigcup_{1\leq l\leq m}\mathcal{D}^{\langle l\rangle})^{+}\alpha:\alpha\in\{-1,1\}\},\qquad(m\in\mathbb{N}).$
We obtain a subshift $X\subset\Sigma^{\mathbb{Z}}$ by excluding from $X_{\mathcal{C}}$ the words in
$\bigcup_{k\in\mathbb{N},\alpha\in\{-1,1\}}\{c^{\langle k,\alpha^{-}\rangle}c^{\langle k+m,\alpha^{+}\rangle},c^{\langle k,\alpha^{-}\rangle}c^{\langle k+m,\alpha^{+}\rangle}:m>1,\ \alpha^{-},\alpha^{+}\in\{-1,1\}\}.$
and the words in $\bigcup_{m\in\mathbb{N}}\mathcal{D}^{\langle m\rangle}.$
Remark: For a word $a\in\mathcal{L}(X)$ that does not have a word in $\mathcal{C}$ as a subword one can find a word $b\in\mathcal{L}(X)$ that has the word $a$ as a prefix and that has a word in $\mathcal{C}$ as a subword: We can distinguish four cases.
For $\alpha\in\{-1,1\},k\in\mathbb{Z}_{+},$ and
$a=g^{\langle+,k,\alpha\rangle},$
set $b$ equal to
$g^{\langle+,k,\alpha\rangle}g^{\langle-,0,\alpha\rangle}=c^{\langle k,\alpha^{-}\rangle}.$
For $(\alpha\in\{-1,1\},k\in\mathbb{N}),$ and
$a=g^{\langle-,k,\alpha\rangle},$
set $b$ equal to
$g^{\langle-,k,\alpha\rangle}c^{\langle k+1,\alpha\rangle},$
which is a subword of
$g^{\langle+,0,\alpha\rangle}g^{\langle-,k,\alpha\rangle}c^{\langle k+1,\alpha\rangle}=c^{\langle k,\alpha\rangle}c^{\langle k+1,\alpha\rangle}.$
For $\alpha\in\{-1,1\},k^{+}>k^{-},$ and
$a=g^{\langle-,k^{-},\alpha^{-}\rangle}g^{\langle+,k^{+},\alpha^{+}\rangle},$
set $b$ equal to
$ag^{\langle-,0,\alpha^{+}\rangle}=g^{\langle-,k^{.},\alpha^{-}\rangle}c^{\langle k+,\alpha^{+}\rangle}$
which is a subword of
$g^{\langle+,k^{+}.k^{-}-1,\alpha^{-}\rangle}g^{\langle-,k^{-}\alpha^{-}\rangle}c^{\langle k^{+},\alpha^{+}\rangle}=c^{\langle k^{+}-1,\alpha^{-}\rangle}c^{\langle k^{+},\alpha^{+}\rangle}.$
For $\alpha\in\{-1,1\},k^{+}\leq k^{-}$, and
$a=g^{\langle-,k^{-},\alpha^{-}\rangle}g^{\langle-,k^{+},\alpha^{+}\rangle},$
set $b$ equal to
$ag^{\langle-,k^{-}-k^{+}+1,\alpha^{+}\rangle}=g^{\langle-,k^{-},\alpha^{-}\rangle}c^{\langle k^{-}+1,\alpha^{+}\rangle}$
which is a subword of
$g^{\langle+,0,\alpha^{-}\rangle}g^{\langle-,k^{-},\alpha^{-}\rangle}c^{\langle k^{-}+1,\alpha^{+}\rangle}=c^{\langle k^{-},\alpha^{-}\rangle}c^{\langle k^{-}+1,\alpha^{+}\rangle}.$
Lemma 9.
Let $R\in\mathbb{N}$, and
$k_{r}\in\mathbb{N},\alpha_{r}\in\{-1,1\},\qquad(1\leq r\leq R).$
such that
$a=\prod_{1\leq r\leq R}c^{\langle k_{r},\alpha_{r}\rangle}\in\mathcal{L}(X).$
Assume that
$Q=\max\{k_{r}:1\leq r\leq R\}-k_{R}>0.$
Then there exist
$\alpha_{q}\in\{-1,1\},\qquad(1\leq q\leq Q),$
such that
$a\prod_{1\leq q\leq Q}c^{\langle k_{R}+q,\alpha_{q}\rangle}\in\mathcal{L}(X).$
Proof.   Set
$r_{1}=\max\{r\in[1,R):k_{r}=k_{R}+1\},$
$r_{q+1}=\max\{r\in[1,r_{q}):k_{r}=k_{q}+1\},\qquad(1\leq q<Q),$
and set
$\alpha_{q}=\alpha^{-}_{r_{q}},\qquad(1\leq q\leq Q).$
$\Box$
Lemma 10.
Let $k\in\mathbb{Z}_{+},\alpha\in\{-1,1\}$. Also let $R\in\mathbb{N}$,
$k_{r}\in\mathbb{N},\alpha_{r}\in\{-1,1\},\qquad(1\leq r\leq R),$
such that
$k_{R}\geq k_{r},\qquad(1\leq r<R).$
(a) Assume that
$a=\prod_{1\leq r\leq R}c^{\langle k_{r},\alpha_{r}\rangle}\in\mathcal{L}(X).$
Then the word $a$ is not semisynchronizing.
(b) Assume that
$a=g^{\langle-,k,\alpha\rangle}\prod_{1\leq r\leq R}c^{\langle k_{r},\alpha_{r}\rangle}\in\mathcal{L}(X)..$
Then the word $a$ is not semisynchronizing.
Proof.   (a) Let
$x^{\langle-\rangle}\in\omega^{\langle-\rangle}(a)$
be transitive. Let $i_{\circ}$ be the maximal index in $(-\infty,0]$ such that
$x^{\langle-\rangle}_{(i-k_{R}-3,i]}\in\{c^{\langle k_{R}+1,\alpha\rangle}:\alpha\in\{-1,1\}\}.$
Then
$y^{\langle+\rangle}=x^{\langle-\rangle}_{(i_{\circ}-k_{R}-3,i_{\circ}]}\prod_{1<q<\infty}c^{\langle k_{R}+q,1\rangle}\in\Gamma^{\langle+\rangle}(a),$
but
$(x^{\langle-\rangle},a,y^{\langle+\rangle})\not\in X.$
(b) Let
$x^{\langle-\rangle}\in\omega^{\langle-\rangle}(a)$
be transitive. In the case that $x^{\langle-\rangle}$ has the suffix
$g^{\langle+,k_{R}-k+1,\alpha\rangle}$
let $i_{\circ}$ be the maximal index in $(-\infty,0]$ such that
$x^{\langle-\rangle}_{(i-k_{R}-4,i]}\in\{c^{\langle k_{R}+2,\alpha\rangle}:\alpha\in\{-1,1\}\}.$
Then
$y^{\langle+\rangle}=c^{\langle k_{R}+1,\alpha^{-1}\rangle}x^{\langle-\rangle}_{(i-k_{R}-4,i]}\prod_{2<q<\infty}c^{\langle k_{R}+q,1\rangle}\in\Gamma^{\langle+\rangle}(a)\in\Gamma^{\langle+\rangle}(a),$
but
$(x^{\langle-\rangle},a,y^{\langle+\rangle})\not\in X.$
Otherwise let $i_{\circ}$ be the maximal index $i$ in $(-\infty,0]$ such that
$x^{\langle-\rangle}_{(i-k_{R}-3,i]}\in\{c^{\langle k_{R}+1,\alpha\rangle}:\alpha\in\{-1,1\}\}.$
Then
$y^{\langle+\rangle}=c^{\langle k_{R}+1,x^{\langle-\rangle}_{i_{\circ}}\rangle}\prod_{1<q<\infty}c^{\langle k_{R}+q,1\rangle}\in\Gamma^{\langle+\rangle}(a),$
but
$(x^{\langle-\rangle},a,y^{\langle+\rangle})\not\in X.$
$\Box$
Proposition 11.
The subshift $X$ is not semisynchronizing
Proof.   The subshift $X$ cannot have a semisynchronizing word. This follows by Lemma 1 from the remark by an application of Lemma 9 and of Lemma 10. $\Box$
The construction of the subshift $X$ is time-symmetric. It follows that the subshift $X$ is not hyposynchronizing.